The market for a product has inverse demand and supply functions given by

p = 290 - 2Qd and p = 10 + 1.5Qs

  1. In what form are these functions in? (2pts)
Answer #1

The form of these functions are Inverse demand and song Inverse demand : P = 290-20d Shverse Subliny: p= 10 +1.5Qs At equilib